OpenStudy (anonymous):

Find the middle term of (x-2)^6

OpenStudy (anonymous):

Its not there, so it should be zero.

OpenStudy (cwrw238):

this will be the 4the term 6C3 x^3 (-2)^3

OpenStudy (cwrw238):

ddo you know how to calculate 6C3 (the number of combination of 3 in 6)?

OpenStudy (anonymous):

Thankss for the help. ☺️☺️

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

The number of terms in the expansion will be 1 more than the power of binomial. In this case, the power is 6, and 6+1=7 so we will have 7 terms all together. To find the 'middle term' of 7, do (7+1)/2 which equals 4. So we need to find the 4th term this binomial! To do this we can use the Binomial Theroem: To find any term in:\[(a+b)^{n}\] we can use: \[t _{k+1}=_{n}C _{k}a ^{n-k}b ^{k}\] we want \[t _{k+1}=t _{4}\] because we are finding the 4th term, so k must equal 3. and n will be the power of the binomial What we have now is: n=6 k=3 a=x b=-2 Lets sub it into the formula \[t _{k+1}=_{n}C _{k}a ^{n-k}b ^{k}\] \[t _{(3)+1}=_{6}C _{3}x ^{(6-3)}(-2)^{3}=20x ^{3}(-8)=-160x ^{3}\]